OpenStudy (anonymous):

select the number of type of solutions for this equation √x^2-28-1=x a. no real roots b. one extraneous root c. two different real roots d. a double root e. three real roots

OpenStudy (anonymous):

b, i think

OpenStudy (anonymous):

x^2=x^2 +58x +29*29, then 0=58x +29^2 ,so x=-29*29/29*2=-29/2